OpenStudy (anonymous):

helppp! Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (mathstudent55):

The point slope formula will help: y - y1 = m(x - x1) where m = slope and the point is (x1, y1)

OpenStudy (anonymous):

yes i know that.. but i just have trouble with solving fractions

OpenStudy (mathstudent55):

OK, here we go: